# VITA: Vision-to-Action Flow Matching Policy

### v2

@@ 5.3 Ablation of Flow Latent Decoding
We investigate the importance of flow latent decoding ($\mathcal{L}_{\text{FLD}}$) for effectively training VITA policies. During training, the decoder is trained by reconstructing from encoder-based latent actions ${\bm{z}}_{1}$, while during inference, the decoder needs to reconstruct from ODE-generated latent actions $\hat{\bm{z}}_{1}$. As a result, the generated latent actions at inference may not decode into meaningful actions, as demonstrated by  Figure˜6. To bridge this gap, we propose two objectives to minimize the discrepancy between encoder-based latents ${\bm{z}}_{1}$ and ODE-generated latents $\hat{\bm{z}}_{1}$: (1) FLD, which decodes $\hat{\bm{z}}_{1}$ into the action space and computes the reconstruction error against the ground-truth action; and (2) FLC, which directly aligns the predicted latent action with the encoded latent target $\bm{z}_{1}=\mathcal{E}_{a}(A)$ using $\|\hat{\bm{z}}_{1}-\bm{z}_{1}\|$.
@@
As shown in Figure˜6, without flow latent decoding (i.e., $\lambda_{\text{FLD}}=0$), the model completely fails to learn a reasonable policy. In contrast, applying FLD succeeds in learning the policy. FLC provides a weaker signal compared to FLD (because FLD directly anchors using the ground-truth actions), resulting in slightly slower convergence. However, we found that using a combination of both objectives yields the best performance due to richer learning signals on both action space and latent space. In short, FLD and FLC are crucial for learning VITA, and in practice a combination of both can achieve optimal results.
@@ 5.4 VITA Denoising Process
Figure 7: Conventional flow matching (a) flows from random Gaussian noise, and gradually “denoise" into raw action chunks by solving ODEs; in contrast, VITA (b) flows from latent images to latent actions, and the latent actions are then decoded to action chunks.
Figure˜7 compares the denoising processes of VITA and conventional flow matching. By jointly optimizing the vision-to-action flow and an action autoencoder with FLD, VITA induces highly structured latent spaces for both vision and action representations. As a result, even the latent image can be directly decoded into a coherent trajectory using the action decoder, which is then progressively refined by the VITA flow matching ODE.
@@ 6 Conclusion
In this work, we developed VITA, an efficient and high-performing visuomotor policy. VITA generates actions in a noise-free manner and directly evolves latent visual representations into latent actions. By leveraging latent image representations as the flow’s source distribution, VITA eliminates the need for complex conditioning mechanisms such as cross-attention, significantly simplifying model architecture and enhancing efficiency. VITA employs vector representations for both latent images and actions, thereby enabling a simple MLP-only architecture for vision-to-action flow. Two key designs are critical to VITA’s success — structuring the latent action space using an action autoencoder and enabling backpropagation of flow latent decoding losses across ODE solving steps — allowing VITA to effectively learn latent action spaces with flow matching in a fully end-to-end manner. Extensive experiments demonstrate that VITA achieves state-of-the-art performance on ALOHA and Robomimic in both real and simulation tasks. VITA achieves $1.5{\times}$-$2.3{\times}$ faster inference compared to conventional flow matching using different conditioning mechanisms.
@@ B.1 Dimensionality Matching for Vision-to-Action Flow
Autoencoder Loss Selection.
@@
Sampling Stochasticity.
We examined VITA variants with different sources of sampling stochasticity. We found that introducing dropout in the network, or variance ($\sigma$) in flow matching (where $\sigma$ injects Gaussian noise into the interpolation path between source and target latents as in OT-CFM) degrades performance. In contrast, adding covariance to the source distribution yields results comparable to the deterministic implementation. Similarly, using a variational objective in the action autoencoder performs on par, whereas applying a variational objective to the image encoder significantly hurts performance (see  Figure˜11).
Appendix C Experiment Results
VITA exhibits efficient and stable learning on AV-ALOHA and PushTFigure˜8) and Robomimic tasks (Figure˜9).
@@ Appendix C Experiment Results
Appendix D Extended Related Works
@@
Imitation Learning for Visuomotor Policy. Imitation learning enables robots to learn complex behaviors by mimicking expert demonstrations. Behavioral cloning, a prominent imitation learning paradigm, frames this as a supervised learning problem, learning a policy that maps observations to actions [26, 27]. Recent advancements in behavioral cloning have widely adopted two modeling methods: generative modeling and autoregressive modeling. Generative methods learn a conditional distribution of actions given an observation. This category includes policies based on conditional variational autoencoders (CVAEs) [26, 27], as well as diffusion [20, 10] and flow matching [30, 14]. On the other hand, autoregressive methods tokenize actions and frame policy learning as a sequence modeling task. These methods predict action tokens sequentially, using next-token prediction [28], next-scale prediction [29], or bi-directional prediction [13]. Generative models ubiquitously require extra conditioning modules (e.g., cross-attention [20], AdaLN [20], FiLM [31, 10]) to inject observations at each step of the generation process. Furthermore, generative and autoregressive methods commonly employ large, expressive networks such as U-Nets or transformers to succeed on complex, high-dimensional robotics tasks. VITA reduces these complexity by formulating the policy as a noise-free and conditioning-free vision-to-action flow; by using vector representations for vision and action, VITA also enables simple and lightweight MLP-only network for policy parameterization, even for challenging bi-manual manipulation tasks.
@@
Diffusion and Flow matching for Generative Modeling. Diffusion [40], grounded in stochastic differential equations (SDEs), generates complex data distributions by sampling from a simple source distribution (typically Gaussian) and iteratively denoising it to the target distribution [19]. Flow matching [18, 41] has been proposed to enable faster training and sampling [41, 32, 42, 18] by modeling the map between source and target distributions with an ordinary differential equation (ODE). Both diffusion and flow matching models have shown strong performance across diverse generative tasks, such as image generation [1, 2, 3, 6, 43, 44], video generation [7, 8], and visuomotor policies [10, 20, 17, 45]. Unlike diffusion, flow matching theoretically places no constraints on the choice of source distribution [32], and a few works have explored leveraging this property to learn the direct transport within the same modality [33, 34], e.g., for image-to-image generation tasks [46, 47]. Recently,  [4] and  [5] extended this to more challenging cross-modal generation between text and image. VITA focuses on learning to bridge vision and action for visuomotor control, where the target modality, action, features sparser data, and lacks semantic structures, compared to text or images, presenting unique challenges. Different from flow matching for image generation, which typically pre-trains and freezes the image encoder and decoder when learning flow matching or diffusion models for image generation [1, 5, 4], VITA resorts to a fully end-to-end pipeline training to effectively learn the latent action space from limited and sparse action data. Furthermore, to enable effective joint training of flow matching and target latent spaces, we propose flow latent decoding to backpropagate action reconstruction losses through the the latent action generation process (ODE solving steps) during training.
Appendix E Discussion of Training Objectives
E.1 Contrastive Latent Alignment
Section˜4.4 introduced two key objectives, FLD and its surrogate FLC, which are both designed to prevent latent space collapse and enable effective VITA learning. Empirically, we find that FLD alone outperforms FLC alone, as it provides a more direct training signal by anchoring the generation process to the ground-truth actions. However, the most robust performance is achieved by combining them, which enforces consistency in both latent space and action space.
Additionally, motivated by the ability of contrastive learning to improve representations and prevent latent collapse [4], we introduced a contrastive loss between vision and action latents [48]. We show that this objective can further boost performance beyond FLD and FLC on some tasks by encouraging the model to learn representations where the similarity between corresponding vision-action pairs is maximized, while the similarity between non-corresponding pairs is minimized.
We employ InfoNCE (Noise-Contrastive Estimation)  [49] for contrastive learning between vision and action. For a given batch of size $N$, the vision latent $\bm{z}_{0,i}$ and action latent $\bm{z}_{1,i}$ from the same data sample are treated as a positive pair. All other non-corresponding combinations $(\bm{z}_{0,i},\bm{z}_{1,j})$ where $i\neq j$ are considered negative pairs. The loss aims to maximize the similarity of positive pairs while minimizing the similarity of negative pairs. The symmetric InfoNCE loss is defined as:
As shown in Figure˜10, the contrastive objective, when used alone, is insufficient for effective VITA learning and underperforms VITA with only the FLD loss (see Figure˜6). However, it provides additional performance gains when combined with the FLD and FLC losses, as it helps create a more robust and well-structured latent space.
E.2 Ablation of the Action Autoencoder
In our experiments, a deterministic autoencoder (AE) learns the target latent space for actions. To investigate the effect of imposing a prior on this latent space, we conducted an ablation study replacing the AE with a variational autoencoder (VAE) [50]. This change introduces a KL divergence regularization term, weighted by $\lambda_{\mathrm{KL}}$, which encourages the encoder’s posterior output, $q(\bm{z}_{1}\!\mid\!A)$, to match a standard normal prior:
As shown by the training curves in Figure˜11(a) and  Figure˜12, incorporating this variational objective with various weights ($\lambda_{\mathrm{KL}}$) can degrade model performance.
E.3 Ablation of the Vision Encoder
As described in our main architecture (Figure˜2), the source latent variable $\bm{z}_{0}=\mathcal{E}_{v}(O)\in\mathbb{R}^{D_{\text{latent}}}$ is produced by a deterministic vision encoder. We performed a similar ablation to assess the impact of a prior on this visual latent space. Specifically, we replaced the deterministic encoder with a variational one that models the posterior $q(\bm{z}_{0}\!\mid\!O)$ and introduces a corresponding KL divergence loss (weighted by $\lambda_{\mathrm{KL}}^{\text{obs}}$):
As shown in Figure˜11(b), using a VAE for the vision encoder drastically degrades performance.
Appendix F Task Settings

### v3

@@ 5.3 Ablation of Flow Latent Decoding
We investigate the importance of FLD ($\mathcal{L}_{\text{FLD}}$) for effectively training VITA policies. As demonstrated by  Figure˜6(a), the generated latent actions at inference may not decode into meaningful actions because of the training-inference gap between encoder-based latent actions ${\bm{z}}_{1}$ and ODE-generated latent actions $\hat{\bm{z}}_{1}$ discussed in  Section˜4.3. To bridge this gap, we propose two objectives that backpropagate through ODEs to explicitly minimize the discrepancy: 1) FLD, which enforces accurate reconstruction in the raw action space by comparing $\mathcal{D}_{a}(\hat{\bm{z}}_{1})$ against the ground-truth action; and 2) FLC, which directly aligns ODE-generated latents $\hat{\bm{z}}_{1}$ with encoder-based latents $\bm{z}_{1}$.
@@
Figure˜6 shows that the model completely fails to learn without FLD (i.e., $\lambda_{\text{FLD}}=0$) due to latent collapse, while applying FLD succeeds in learning the policy. FLC provides a weaker signal compared to FLD (because FLD directly anchors generation using the ground-truth actions), resulting in slightly slower convergence. However, we found that using a combination of both objectives in practice yields the best performance due to richer learning signals in both raw and latent action spaces.
@@ 5.4 VITA Denoising Process
Figure˜7 compares the denoising processes of VITA and conventional FM. By jointly optimizing the vision-to-action flow and the latent action space via FLD, VITA induces highly structured latent spaces for both vision and actions. Consequently, even latent images express action semantics: latent images can be directly decoded into coherent action trajectories using the action decoder and are then progressively refined by the ODE.
Figure 7: Conventional flow matching (a) evolves random Gaussian noise into action chunks by solving ODEs, gradually denoising noisy actions; in contrast, VITA (b) flows directly from latent images, and refines latent actions over ODE steps. Notably, by bridging vision and actions via VITA learning, the latent visual representations begin to express action semantics and can be decoded into action trajectories using the action decoder.
@@ 6 Conclusion
We developed VITA, an efficient and high-performing visuomotor policy, which generates actions in a noise-free manner. VITA directly evolves visual latents into latent actions. VITA eliminates the need for conditioning modules such as cross-attention, simplifying architectures and enhancing efficiency. When employing vector representations for both latent images and actions, VITA reduces the flow matching network to a conditioning-free vector-to-vector mapping, enabling a simple MLP-only architecture for complex visuomotor tasks. Two designs are critical to VITA’s success: learning a structured latent action space as the flow matching target, and backpropagating across ODE solving steps, allowing VITA to effectively learn in a fully end-to-end manner. Extensive experiments demonstrate that VITA achieves state-of-the-art performance on ALOHA and Robomimic in both real and simulation tasks. VITA achieves $1.5{\times}$-$2{\times}$ faster inference and 18.6%-28.7% less memory usage compared to conventional flow matching.
Appendix
@@ B.1 Dimensionality Matching for Vision-to-Action Flow
B.2 Ablation of Frozen Target for Flow Matching
Learning a flow when the target distribution lies in a learned latent space is inherently challenging. Jointly optimizing the flow and the latent encoder creates a moving target, as the latent space continually shifts during training; it also introduces a training-inference gap: flow matching does not guarantee that ODE-generated latents are decodable during inference, since the decoder is trained to reconstruct encoder-based latents.
Figure 8: Success rates (left) and log-scaled action MSEs (right) comparing end-to-end VITA training with VITA using a frozen action AE on ThreadNeedle. EMA = 0.9.
A natural solution is to follow the strategy in latent diffusion models [stablediffusion]: pre-train the latent space on large-scale data and freeze it when training the flow. We evaluate this approach by pre-training the action autoencoder for 100k steps using a reconstruction loss, $\mathcal{L}_{\text{AE}}$, then freezing it while training the VITA policy for 25k steps (same as VITA trained from scratch). As shown in Figure˜8, this frozen-latent setup yields poor online success rates and high offline MSEs that plateau early. In contrast, end-to-end VITA training enabled by FLD performs substantially better. The underlying issue is that, unlike image generation, robotics action data is sparse and limited; pure latent pre-training produces a weak representation that cannot be improved once the latent space is frozen.
B.3 Contrastive Latent Alignment
Section˜4.4 introduced two key objectives, FLD and its surrogate FLC, which are both designed to prevent latent space collapse and enable effective VITA learning. Empirically, we find that FLD alone outperforms FLC alone, as it provides a more direct training signal by anchoring the generation process to the ground-truth actions in the raw action space. However, the most robust performance is achieved by combining them, which enforces consistency in both latent space and action space.
Additionally, motivated by the ability of contrastive learning to improve representations and prevent latent collapse [crossflow], we introduced a contrastive loss between vision and action latents [clip-openai]. We show that this objective can further boost performance beyond FLD and FLC on some tasks by encouraging the model to learn representations where the similarity between corresponding vision-action pairs is maximized, while the similarity between non-corresponding pairs is minimized.
We employ InfoNCE (Noise-Contrastive Estimation) for contrastive learning between vision and action. For a given batch of size $N$, the vision latent $\bm{z}_{0,i}$ and action latent $\bm{z}_{1,i}$ from the same data sample are treated as a positive pair. All other non-corresponding combinations $(\bm{z}_{0,i},\bm{z}_{1,j})$ where $i\neq j$ are considered negative pairs. The loss aims to maximize the similarity of positive pairs while minimizing the similarity of negative pairs. The symmetric InfoNCE loss is defined as:
As shown in Figure˜9, the contrastive objective, when used alone, is insufficient for effective VITA learning and underperforms VITA with only the FLD loss (see Figure˜6). However, it provides additional performance gains when combined with the FLD and FLC losses, as it helps create a more robust and well-structured latent space.
B.4 Ablation of the Variational Action Autoencoder
In our experiments, a deterministic autoencoder (AE) learns the target latent space for actions. To investigate the effect of imposing a prior on this latent space, we conducted an ablation study replacing the AE with a variational autoencoder (VAE) [vae]. This change introduces a KL divergence regularization term, weighted by $\lambda_{\mathrm{KL}}$, which encourages the encoder’s posterior output, $q(\bm{z}_{1}\!\mid\!A)$, to match a standard normal prior:
As shown by the training curves in Figure˜10(a) and Figure˜21, incorporating this variational objective with various weights ($\lambda_{\mathrm{KL}}$) can degrade model performance.
B.5 Ablation of the Variational Vision Encoder
As described in our main architecture (Figure˜2), the source latent variable $\bm{z}_{0}=\mathcal{E}_{v}(O)\in\mathbb{R}^{D_{\text{latent}}}$ is produced by a deterministic vision encoder. We performed a similar ablation to assess the impact of making the vision encoding stochastic. Specifically, we replaced the deterministic encoder with a variational one that models the posterior $q(\bm{z}_{0}\!\mid\!O)$ and introduces a corresponding KL divergence loss (weighted by $\lambda_{\mathrm{KL}}^{\text{obs}}$):
As shown in Figure˜10(b), using a VAE for the vision encoder drastically degrades performance.
Although using a VAE to model the source distribution introduces stochasticity that is often desirable for generative tasks emphasizing diversity [crossflow], we found that making the visual latent encoding stochastic degrades VITA performance. In robotics tasks that demand extremely high precision, such as ThreadNeedle, where millimeter-level errors cause complete failure, variational objectives tend to blur latent representations, discarding critical visual details. As a result, deterministic vision encodings yield substantially better precision and performance.
B.6 Ablation of Network Architecture
VITA, when using vector representations for both vision and actions, reduces the flow matching network to a conditioning-free vector-to-vector mapping. We find that even an MLP-only architecture can successfully learn challenging, high-precision visuomotor tasks, including bimanual manipulation with active vision on AV-ALOHA. To the best of our knowledge, VITA is the first visuomotor policy to master such complex tasks using MLPs.
B.6.1 FM using MLP
Figure 11: Reward curves and the action MSEs (log-scaled) of MLP-only VITA and MLP-only FM on PushT. FM learning is ineffective because lack of precision, and performs poorly online, necessitating effective architectures to process action chunks and fuse in conditions.
We evaluate FM using the same 4-layer MLP architecture as VITA (with a similar parameter count of $\sim$30M). Concretely, we remove self-attention from the DiT-based FM network and retain only AdaLN (which is also MLP-based) for visual conditioning. However, the MLP-only FM fails to learn effective policies: on PushT, the reward remains around 0.4 and success rate remains 0% even after 100K training steps, as the task requires high-precision control (successful only when reward exceeds 0.95). In contrast, VITA and transformer-based FM achieve 88% and 83% success, respectively. This failure arises because an MLP is insufficient for processing noisy action chunks and integrating visual conditioning at each denoising step, leading to poor control precision, as shown in Figure˜11(b), where action MSEs of MLP-based FM plateau while VITA yields significantly lower MSEs. Since success rates are sparse on PushT, we also report online reward curves (see Figure˜11(a)) to more clearly compare MLP-based FM and VITA.
B.6.2 VITA using Transformers
Additionally, we show that VITA is not limited to vector-based features or MLP. We evaluate VITA using grid-based features (in particular, $9\times 512$ spatial features obtained via ResNet-18), and use transformer for the flow matching network. The flow matching network does not require any conditioning for spatial tokens such as costly cross-attention compared to FM using transformers.
We evaluate VITA on multiple challenging tasks, demonstrating that VITA yield high success rates (see Figure˜12) while retaining the efficiency gains (see Table˜1 and Appendix B.7).
Figure 12: Comparing success rates of VITA using MLP and vector-based features and VITA using transformers and grid-based features.
B.6.3 Latent Collapse
Our motivation for FLD stems from two key observations: first, the multi-stage pipeline inspired by latent diffusion utilizing a pre-trained autoencoder is suboptimal for action data (Section˜B.2); and second, joint training of the flow and target latent space leads to latent collapse (Section˜5.3).
Notably, very recent work  [dsd-e2e-diffusion] resonates with these findings, and independently identified a similar failure mode in latent diffusion for image generation. They observed that the diffusion objective can undermine representation quality during end-to-end training, and proposed Diffusion as Self-Distillation (DSD) [dsd-e2e-diffusion] to stabilize latent representation learning. While FLD and DSD employ different approaches in distinct domains, both methods corroborate the necessity of end-to-end optimization while stabilizing latent representation learning.
B.7 Extended Efficiency Metrics
B.7.1 Training Time & Space Overheads of Flow Latent Decoding
Recall that we introduce FLD during training to enable effective end-to-end learning, which requires solving 6-step ODEs. Although this inevitably adds some training-time overhead, we trade a modest increase in training cost for significantly more efficient inference, which is critical for real-time robotic deployment. For MLP-based VITA, the training time increases from 0.677 ms/chunk (without FLD) to 0.740 ms/chunk (with FLD), corresponding to only a 9.3% overhead. For transformer-based VITA, the cost rises from 0.766 ms/chunk to 0.951 ms/chunk, a 24.1% overhead. A similar trend appears for GPU memory usage: for MLP-based VITA, peak training memory increases from 2716.62 MiB (without FLD) to 2835.86 MiB (with FLD), a relatively small 4.4% increase. Despite these modest training-time and memory overheads, FLD enables stable end-to-end optimization and yields markedly improved inference-time efficiency, a critical requirement for real-time robotic deployment. Additionally, VITA maintains the lowest training-time memory usage compared to all baselines even with FLD enabled, while achieving comparable training latency.
B.7.2 Training Memory Usage
We have reported the absolute peak memory usage of each policy architecture in Table 1, measured via torch.cuda.max_memory_allocated for fair comparison. Importantly, we exclude the shared visual encoder, which transforms raw RGB images into visual latents and is used identically across all methods. This observer accounts for approximately 1732.5 MiB of peak GPU memory across all methods, and does not help differentiating memory usage across policies. For instance, after removing observer cost, VITA (MLP) peaks at 333.86 MiB, FM (MLP, AdaLN) peaks at 413.95 MiB, and FM (Transformer, Cross-Attn) reaches 529.16 MiB. These numbers reflect the true architectural and conditioning differences in the policy modules themselves. Consequently, VITA (MLP) reduces peak memory usage by 19.4% relative to FM (MLP, AdaLN), 18.6% relative to FM (Transformer, AdaLN), and 36.9% relative to FM (Transformer, Cross-Attn), highlighting its inference-time efficiency gains from a conditioning-free formulation.
In Table˜6, we report the absolute peak memory usage during training for each policy. As with inference, we ablate the memory consumed by the shared visual encoder, and record peak memory after vision encoding. For VITA, we also include the peak usage observed during the flow_latent_decoding phase, which accounts for the highest memory load. This choice ensures that the reported number reflects the true upper bound of VITA training footprint with FLD.
Table 6: Comparison of the conditioning parameter overhead, training-time cost, and training-time memory usage of VITA and baselines, grouped by the type of visual latents used (“Vector” or “Grid” based). Metrics include (i) parameters introduced solely by conditioning modules, (ii) training time per chunk (ms), and (iii) peak GPU memory during training (MiB).
Visual Model Architecture Conditioning Cond. Params (M) Time Memory
Vector VITA MLP N/A 0.00 0.740 2835.86
FM MLP AdaLN 11.82 0.664 2926.60
FM Transformer AdaLN 6.58 0.697 3071.88
FM U-Net FiLM 11.33 0.782 3676.38
DDPM U-Net FiLM 9.49 0.779 3643.04
Grid VITA Transformer N/A 0.00 0.951 2977.10
FM Transformer Cross-Attn 4.47 0.812 3585.06
B.8 Control Precision and Sampling Stochasticity
B.8.1 VITA with Sampling Stochasticity
Since stochastic image encodings degraded performance (as discussed in Section˜B.5), and we hypothesized this was due to precision loss from blurred latent representations, we further investigated the trade-off between stochasticity and control performance. We examined VITA variants with different sources of sampling stochasticity. Introducing dropout in the network, or variance ($\sigma$) in flow matching, where $\sigma$ injects Gaussian noise along the interpolation path, consistently reduced performance. In contrast, adding covariance to the source distribution produced results comparable to deterministic encoding. Likewise, using a variational objective in the action autoencoder performed similarly to the deterministic action autoencoder, whereas applying a variational objective to the image encoder significantly harmed performance (see Figure˜10).
These findings resonate with our observation that DP underperforms FM or VITA on ALOHA tasks that require high precision. DP is based on an SDE (stochastic differential equation) while FM uses a deterministic ODE formulation. FM introduces stochasticity only through sampling the Gaussian prior, VITA goes even further by removing the Gaussian prior sampling, instead using a visually grounded and deterministic initial state for the flow. Together, these results can suggest a broader trend: for fast and precise real-time control required by many tasks, reducing stochasticity can be beneficial in, e.g., speeding up convergence, producing more precise and faster policies.
B.8.2 Analyzing Under-Performance of DP and ACT on ALOHA Tasks
We followed DP and ACT implementation from LeRobot [cadene2024lerobot]. However, DP and ACT perform poorly ( 40% SRs compared to 80%-90% of VITA and FM) on ALOHA tasks such as ThreadNeedle and PourTestTube, which demand high precision.
This section examines the stringent success criteria of these tasks and explains why even small action errors lead to failures. We further show that VITA and FM achieve substantially better action precision with far fewer training steps, which contributes to their superior performance on these precision-demanding tasks.
Success Criteria. We observed that DP and ACT learn reasonable trajectories, but small millimeter-level errors lead to binary failures. For example, because success requires completing all five subtasks on ThreadNeedle, failing at the third stage still counts as a full failure, yielding low SRs when the average reward exceeds 3. In contrast, as depicted in Figure˜13, both VITA and FM learn sufficiently precise control to complete the final subtask with fewer training steps, achieving high success rates on these tasks.
Figure 13: Success rate and reward curves of VITA and DP on ThreadNeedle.
As shown in Figure˜14, DP can complete multiple subtasks yet fail the episode due to millimeter-level precision errors, such as threading a needle through a very small opening. In contrast, we find that VITA and FM exhibit higher control precision on these tasks.
Figure 14: A failure and a success case on ThreadNeedle. DP may complete most subtasks but still fail the final insertion due to millimeter-level errors.
Action Precision. We now examine why VITA and FM overall achieve higher action precision. As shown in Table˜3, VITA outperforms all baselines on most tasks, and FM achieves SRs comparable to VITA. To better understand this, we analyze the offline action MSE during training for VITA, FM, DP, and ACT in Figure˜15. We observe that VITA and FM consistently converge to lower MSEs, whereas ACT plateaus at substantially higher errors, and DP uses much more training steps to converge. This trend aligns with prior findings (e.g., on image generation) showing that flow matching methods enjoy faster convergence [lipman2024flowmatchingguidecode] and can achieve higher generation fidelity [fm-fidelity].
Figure 15: Comparison of action MSE on three real-world single-arm ALOHA tasks.
B.9 Autoencoder Loss Selection.
@@ B.8.2 Analyzing Under-Performance of DP and ACT on ALOHA Tasks
B.10 Success Rate Curves during VITA Learning
VITA exhibits efficient and stable learning on AV-ALOHA and PushT (Figure˜16) and Robomimic tasks (Figure˜17).
@@ B.10 Success Rate Curves during VITA Learning
Appendix C Extended Related Works
@@
Imitation Learning for Visuomotor Policy. Imitation learning enables robots to learn complex behaviors by mimicking expert demonstrations. Behavioral cloning, a prominent imitation learning paradigm, frames this as a supervised learning problem, learning a policy that maps observations to actions [zhao2023learning, lee2024interact]. Recent advancements in behavioral cloning have widely adopted two modeling methods: generative modeling and autoregressive modeling. Generative methods learn a conditional distribution of actions given an observation. This category includes policies based on conditional variational autoencoders (CVAEs) [zhao2023learning, lee2024interact], as well as diffusion [dit-policy, chi2023diffusion] and flow matching [fmp-affordance, fmp-3d-consist]. On the other hand, autoregressive methods tokenize actions and frame policy learning as a sequence modeling task. These methods predict action tokens sequentially, using next-token prediction [fu2024context], next-scale prediction [carp], or bi-directional prediction [su2025dense]. Generative models ubiquitously require extra conditioning modules (e.g., cross-attention [dit-policy], AdaLN [dit-policy], FiLM [perez2018film, chi2023diffusion]) to inject observations at each step of the generation process. Furthermore, generative and autoregressive methods commonly employ large, expressive networks such as U-Nets or transformers to succeed on complex, high-dimensional robotics tasks. VITA reduces these complexity by formulating the policy as a noise-free and conditioning-free vision-to-action flow.
@@
Diffusion and Flow matching for Generative Modeling. Diffusion [ho2020denoising], grounded in stochastic differential equations (SDEs), generates complex data distributions by sampling from a simple source distribution (typically Gaussian) and iteratively denoising it to the target distribution [sohl2015deep]. Flow matching [lipman2023flowmatchinggenerativemodeling, liu2022flowstraightfastlearning] has been proposed to enable faster training and sampling [liu2022flowstraightfastlearning, tong2024bridgesscore, scaling-rectified-flow, lipman2023flowmatchinggenerativemodeling] by modeling the map between source and target distributions with an ordinary differential equation (ODE). Both diffusion and flow matching models have shown strong performance across diverse generative tasks, such as image generation [stablediffusion, dit, sit, conddiffusion, liu2024alleviating, ren2024flowar], video generation [video-diff, li2023videogen], and visuomotor policies [chi2023diffusion, dit-policy, black2410pi0, liu2024rdt]. Unlike diffusion, flow matching theoretically places no constraints on the choice of source distribution [tong2024bridgesscore], and a few works have explored leveraging this property to learn the direct transport within the same modality [i2i-flow-interpolants, i2i-flow-sim-free], e.g., for image-to-image generation tasks [i2i-flow-boosting, i2i-flow-rec]. Recently,  [crossflow] and  [flowtok] extended this to more challenging cross-modal generation between text and image. VITA focuses on learning to bridge vision and action for visuomotor control, where the target modality, action, has sparser data and lacks semantic structures, compared to text or images, presenting unique challenges. Different from flow matching for image generation, which typically pre-trains and freezes the image encoder and decoder when learning flow matching or diffusion models for image generation [stablediffusion, flowtok, crossflow], VITA resorts to a fully end-to-end pipeline training to effectively learn the latent action space from limited and sparse action data. Furthermore, to enable effective joint training of flow matching and target latent spaces, we propose flow latent decoding to backpropagate action reconstruction losses through the the latent action generation process (ODE solving steps) during training.
Appendix D Task Settings
To comprehensively evaluate the effectiveness of VITA across varying levels of difficulty and action dimensionality, we conduct extensive experiments on both single-arm and bimanual manipulation tasks. The action dimensionality spans from 2 to 21, and the tasks include both short- and long-horizon settings. Overall, AV-ALOHA tasks are particularly challenging due to their 21D action spaces, non-stationary observations introduced by the active-vision camera, and long-horizon, precision-demanding task structure (see Figure˜18 for real-world examples). Single-arm ALOHA tasks are also challenging due to randomness, such as varying object types and object poses.
The specifications for each dataset are shown in Table˜7. Following the practice of AV-ALOHA [chuang2024active], we train all policies at 8.33 FPS ($25/3$) for simulated AV-ALOHA tasks and at 11 FPS for real AV-ALOHA tasks, and interpolate to 25 FPS and 33 FPS, respectively, for inference.